OpenStudy (anonymous):

How to show work solving 27.8 x 9.6=

OpenStudy (nurali):

Try and break it down (but never forget about a component) I first break up the 27 from its .8, then keep simplifying into sums that I can do in my head. 27 x 9.6 =27 x 9 + 0.6 x 27 = 25x9 + 2x9 + 0.6x2 + 25x0.6 = 225+ 18 + 1.2 + 15 = 259.2 + .8 x 9.6 = .8 x 9 + .8 x .6 = 7.2 + 0.48 = 7.68 Giving a grand total of 266.88

OpenStudy (nurali):

Welcome to Openstudy.